I'm having trouble with this question about vectors:
A pilot wishes to fly from airport A to airport B. Airport B is located due west from airport A. The pilot's aircraft averages a speed of 160 mi/hr. If the pilot is flying in a wind which blows out of the north at speed of 40 mi/hr, then the pilot should head his plane in the direction of vector...?

Answer by Alan3354(69443)

A pilot wishes to fly from airport A to airport B. Airport B is located due west from airport A. The pilot's aircraft averages a speed of 160 mi/hr. If the pilot is flying in a wind which blows out of the north at speed of 40 mi/hr, then the pilot should head his plane in the direction of vector
=====================
Did you see my other answer?
------
The hypotenuse is 160.
Side opposite = 40
-----------
The groundspeed = sqrt(160^2 - 40^2) = sqrt(24000) = 40sqrt(15)
----
In the x-y plane --> vector -40*sqrt(15)i + 40j